OpenStudy (anonymous):

what is the factored form of y = x^2 – 10x + 25

OpenStudy (igreen):

What two numbers get 25 when they're multiplied and get -10 when they're added?

OpenStudy (anonymous):

5 x 5

OpenStudy (igreen):

No, 5 + 5 = 10

OpenStudy (anonymous):

-5?

OpenStudy (shamim):

-5*(-5)=?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (shamim):

-5-5=?

OpenStudy (anonymous):

0

OpenStudy (igreen):

So we have: (x - 5)(x - 5)

OpenStudy (anonymous):

thats it?

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

i green can you help me my question is at the top iam new here